2) A coin is weighted so that there is a 60% chance of it landing on heads when flipped. The coin is flipped 13 times.

Find the probability that at most 8 of the flips resulted in "heads". Round your answer to 4 decimal places.
Answer #1

X : number of flips that resulted head.

X ~ bin (n,p)

where, n = 13

p = 60% = 0.60

q= (1-p) =(1-0.60)= 0.40

the pmf of the distribution be:-

P(X = z) = ()?q=== (13)0.6+0.48- P(X = 1) = 10.6 * 0.413-1

the probability that at most 8 of the flips resulted in "heads" be:-

P(X <8)

=1-P(X = 9) + P(X = 10) + P(X = 11) + P(X = 12) + P(X = 13)

=1-(5).೮ +0.419-+ (1)೦.೮° +0.43-10+ (1) 0.6 +0.419-11, (13).6³2 +0.413-12 + (3) .6³3 + 0.413-13)

0.6470

*** if you have any doubt regarding the problem please write it in the comment box...if you are satisfied please give me a LIKE if possible...